Question 5 of 15
Multiple Response: Please select all correct answers and click "submit."
Which choices are equivalent to the quotient below? Check all that apply. Square root 50 /square root 10


A. Square root 3
B. 15/3
C. Square root 25 /square root 5
D. 5
E. Square root 5
F. Square root10 /square root 2

Answer by stanbon(75887)

Please select all correct answers and click "submit."
Which choices are equivalent to the quotient below? Check all that apply. Square root 50 /square root 10 = sqrt(50/10) = sqrt(5)
---------------------------
A. Square root 3
B. 15/3
C. Square root 25 /square root 5 = sqrt(5) ------Submit
D. 5----submit
E. Square root 5
F. Square root10 /square root 2 = sqrt(5) -------Submit
--------------------
Cheers,
Stan H.
---------------